OpenStudy (anonymous):

What happens to the speed of an electromagnetic wave as it travels through a vacuum?

OpenStudy (abb0t):

It goes to a constant value called c, which is 300000000m/s \[3.0 \times 10^8 \frac{ m }{ s }\] same as speed of light.

OpenStudy (anonymous):

constant speed c.which is equal to 2.9979246*10^8 m/s. this is from c=1/root of mu*epsilon

OpenStudy (vincent-lyon.fr):

Exact value is 299792458 m/s